A planes flies south at 192mph. Two hours later another plane leaves at 960mph. when will the second plane pass the first plane.
Answer by ewatrrr(24785)


Hi
A planes flies south at 192mph. Two hours later another plane leaves at 960mph
Let x and (x-2)represent times traveled respectively
D = rt (Distance traveled the same for both planes)
192x = 960(x-2)
1920 = 768 x
x = 2.5 hrs , time traveled by slower plane.
Second plane will pass the first plane in 1/2 hr OR 30min.
CHECKING our Answer***
192*2.5 = 480mi = 960*(1/2)